The number of long distance calls between two cities in a certain time period varies jointly with the populations of the cities and inversely with the distance between them. If 25,000 calls are made between two cities 200 miles apart with populations of 50,000 and 20,000, how many calls are made between cities of 75,000 and 40,000 that are 400 miles apart?
Answer by stanbon(75887) 
 The number of long distance calls between two cities in a certain time period varies jointly with the populations of the cities and inversely with the distance between them.
If 25,000 calls are made between two cities 200 miles apart with populations of 50,000 and 20,000,
# calls = k(pop/d)
Find "k":
25000 = k(70000/200)
25000 = k(350)
k = 500/7
---
Therefore the equation for this problem is # calls = (500/7)pop/d
-----------------------------------
how many calls are made between cities of 75,000 and 40,000 that are 400 miles apart?
#calls = (500/7)115000/400
# calls = 20535.71
Rounding down you get # calls = 20,535
================================================
Cheers,
Stan H.
| |
|